OpenStudy (arkay):

log(x+5)-log x=3 answers 0.0505 0.0050 2.5 199.8?

OpenStudy (zzr0ck3r):

\(\log(x+5)-\log (x)=3\\\log(\dfrac{x+5}{x})=3\) Now use \(\log_a(b)=c\) if and only if \(a^c=b\)

OpenStudy (karim728):